OpenStudy (anonymous):

g(x)=3(1-t)^2-4(1-t)+14 g(1-t)=?

OpenStudy (michele_laino):

is g(x) like below: g(x) = 3x^2-4x+14

OpenStudy (michele_laino):

if yes, the replacing x with 1-t, we get: \[\large g\left( {1 - t} \right) = 3{\left( {1 - t} \right)^2} - 4\left( {1 - t} \right) + 14\]

OpenStudy (michele_laino):

then*

OpenStudy (michele_laino):

please simplify the above expression

OpenStudy (anonymous):

I'm having trouble foiling, I think and then simplifying. I got to \[3-6t+3t ^{2}-4+4t+14\]

OpenStudy (freckles):

I'm confused about the question g(x)=3(1-t)^2-4(1-t)+14 means g(x) is constant unless t is an expression of x

OpenStudy (freckles):

if g(x)=a where a is constant then g(1-t)=a also